# Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models

### v1

@@ 4 Results
Figure 3: Classifier selectivity across model layers. Selectivity is defined as the difference between linguistic and control task accuracy, with higher values indicating better generalization rather than memorization. The top row shows lexeme selectivity; the bottom row shows inflection selectivity. Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right). Each line represents a different model.
4.1 Lexical Identity
Linear Regression Classifiers
Lemma prediction accuracy using linear regression shows a consistent declining pattern across layer depth for all models, beginning with accuracy between 0.8-1.0 in early layers and declining toward final layers (Figure 2, top left). DeBERTa-v3-Large exhibits the sharpest decline (0.9 $\rightarrow$ 0.3), while Pythia-6.9B maintains the highest sustained performance in later layers. Performance varies significantly across parts of speech, with models achieving highest accuracy on frequent categories like nouns and verbs, while struggling with infrequent categories like pronouns and prepositions due to limited training examples (Table 10).
MLP Classifiers
Two-layer MLP classifiers show higher overall accuracies and less pronounced decline compared to linear regression (Figure 2, top middle). Pythia models show the strongest performance throughout, while GPT-2, BERT and Qwen2.5 variants show the largest improvements from linear to MLP classifiers. The same frequency-dependent performance patterns persist across parts of speech (Table 11).
Random Forest Classifiers
Random forest results for lemma prediction are not reported due to computational constraints arising from thousands of unique lemma classes.
4.2 Inflectional Morphology
Linear Regression Classifiers
Inflectional morphology prediction using linear regression exhibits consistently high performance (0.9-1.0) across all layers and models (Figure 2, bottom left). The accuracy curves show little degradation from early to late layers. Between classifiers trained on different language model representations, there is minimal variance in accuracy. While most morphological forms achieve high accuracy, comparative and superlative adjective forms are most challenging (Table 8).
MLP Classifiers
MLP classifiers show nearly identical patterns to linear regression, maintaining accuracy between 0.9-1.0 across all layers (Figure 2, bottom middle). This highlights the fact that inflectional morphology is encoded very linearly, as additional nonlinear capacity provides no meaningful benefit. The pattern of lower performance on comparative and superlative forms persists with MLP classifiers (Table 9).
Random Forest Classifiers
Random forest classifiers exhibit lower overall accuracy (0.65-0.90) and a declining trend across layers (Figure 2, bottom right), contrasting sharply with the flat, high-performance curves of linear regression and MLP classifiers.
Figure 4: Fraction of maximum components needed to achieve different explained variance thresholds across layers. Each subplot shows a different variance threshold (50%-100%), with lines colored by model. Note that 100% explained variance always requires the full component set. Models fall into two groups: in one, dimensionality declines gradually; in the other, the middle layers require only one component. See additional plots in Figure 8.
@@ 4.2 Inflectional Morphology
5 Analysis
Our experiments reveal clear patterns in how transformer models encode linguistic information. We organize our analysis around six key aspects: distinct encoding patterns, classifier complexity, generalization versus memorization, representation dimensionality, the role of model architectures & training dynamics, and the effects of tokenization.
Distinct Encoding Patterns
We find consistent and distinct encoding patterns for lexical and inflectional information. Lexical information is most accessible in early layers (0-2) and then declines, while inflectional information remains uniformly high across all layers (Figure 2). This separation suggests that lexical identity is captured as an early, surface-level feature, while morphological features remain explicit throughout. This extends the hierarchical view of transformer representations advanced by Tenney et al. (2019) and Jawahar et al. (2019), by showing that surface morphology is not relegated to early layers alone but is preserved as a linearly decodable signal at every depth.
Differences in Classifier Complexity
We find that different types of linguistic information require different levels of classifier complexity to extract (Figure 2), which we quantify using the linear separability gap from § 2. Inflectional features show a small gap across all models, indicating that they remain linearly decodable. Lexical features, by contrast, show an increasingly large gap in later layers, suggesting that lexical features become progressively more non-linearly encoded (Figure 7 in Appendix C).
Generalization vs. Memorization
A key question is whether classifiers detect genuine linguistic structure or simply memorize training examples. We implement control tasks following Hewitt and Liang (2019), where words are randomly but consistently mapped to arbitrary labels. High classifier selectivity - the accuracy difference between linguistic task accuracy and control task accuracy - indicates generalization over memorization.
As shown in Figure 3, inflection prediction shows strong selectivity (0.4-0.6), especially in later layers, suggesting that models develop generalizable abstractions for morphological patterns. In contrast, lemma prediction has near-zero selectivity across all models, layers, and classifier types, suggesting that lexical information is retrieved primarily through memorization. [^7] This is perhaps unsurprising given the nature of the lemma task: it involves thousands of unique classes, many of which appear infrequently in the training data.
This aligns with the notion that morphological rules are systematic and generalizable (e.g., adding -ed to form past tense), while lexical identity is item-specific. It also connects to previous work on the increasing contextualization of representations in later layers of transformer models (Ethayarajh, 2019). Subramani et al. (2022) similarly found that steering vectors become less effective in later layers and that representations reflect semantic similarity less in later layers.
Taken together, these suggest that abstract linguistic properties become increasingly entangled with contextual information, making them increasingly harder to extract with linear methods but still recoverable through non-linear transformations.
Representation Dimensionality
[…]
Models fall into two groups (Table 2). In one group (BERT, DeBERTa, Gemma, Llama 3.1), dimensionality declines gradually across layers. In the other (GPT-2, Qwen2.5, Pythia, OLMo-2), the middle layers require just one component to reach 50-99% variance, before rising again towards the final layers of the model
Importantly, these low-dimensional middle layers are where inflectional features are most linearly separable and where lexical features show the largest non-linear gap (see Figure 7). In other words, middle layers learn to compress representations into a tight subspace that cleanly encodes inflectional morphology but entangles lexemes. Later layers expand dimensionality again to support generation or deeper contextualization while preserving linearly accessible inflectional information.
Model Architectures
Despite differences in architecture (encoder vs. decoder), model size, and training objectives, all models show remarkably consistent patterns: relatively flat, high-accuracy curves for inflection prediction and declining accuracy for lemma prediction in later layers (Figure 2). This holds across causal decoder (GPT-2, Pythia, OLMo-2, Qwen, Gemma, Llama 3.1) and bidirectional encoder models (BERT, DeBERTa).
Surprisingly, instruction-tuned models like Qwen-2.5-Instruct, Llama-3-8B-Instruct, and Pythia-6.9B-Tulu display nearly identical trends to their base counterparts (Appendix CTable 7). This suggests that instruction tuning does not substantially alter how lexical or inflectional features are encoded.
Training Dynamics
We also examined how these patterns develop during pre-training by training classifiers at intermediate model checkpoints for both OLMo-2 and Pythia (Figure 9 and Figure 10). Lexeme classification reaches peak accuracy early in training before slowly declining, while inflectional accuracy remains high and stable across both checkpoints and layer depth. This result suggests that models establish the separation between lexical and inflectional information early and maintain this structure throughout training. Taken together with our model architecture findings, we find that separating lexical and inflectional information facilitates effective next-token prediction.
Effects of Tokenization
Figure 5: Effect of tokenization strategy on analogy completion rank. Each point corresponds to a model (color) and analogy (shape). The x-axis is the rank using whole-word representations. The y-axis is the rank using tokenized representations. Here, rank means the position of the expected word when all vocabulary words are sorted by similarity to the resulting embedding from vector arithmetic; lower is better. Points above the gray $y{=}x$ line mean tokenization hurts performance.
Tokenization is an essential component of language modeling. To test how tokenization influences our findings, we use analogy completion tasks (e.g., man:king::woman:?) and compare two approaches: averaging subtoken embeddings after standard tokenization and summing embeddings from whole-word tokens.
Whole-word representations markedly outperform averaged subtokens across all models (Figure 5), implying that linguistic regularities are primarily stored in whole-word embeddings rather than compositionally across subtokens. Despite tokenization effects, our classifier results show consistent patterns across models using different tokenizers (see Table 5), indicating robust encoding of lexical and morphological information.
6 Related Work
Probing for linguistic information
Probing studies use supervised classifiers to predict linguistic properties from model representations (Alain and Bengio, 2017; Adi et al., 2017). Methodological improvements include control tasks to ensure probe selectivity (Hewitt and Liang, 2019). Extensive work has established that transformer models learn hierarchical linguistic structures, with different layers specializing in different information types: lower layers capture surface features and morphology, middle layers encode syntax, and upper layers represent semantics and context (Jawahar et al., 2019; Tenney et al., 2019; Liu et al., 2019; Rogers et al., 2020). Most relevant to our work, Vulić et al. (2020) found that lexical information concentrates in lower layers, while Ethayarajh (2019) showed that representations become increasingly context-specific in higher layers.
Activation steering
Beyond probing, recent work has explored manipulating model behavior by intervening on internal representations. This includes steering vectors (Subramani et al., 2022), inference-time interventions (Li et al., 2023), representation editing (Meng et al., 2022), sparse autoencoders for feature discovery (Bricken et al., 2023), and causal mediation analysis (Vig et al., 2020). These methods complement our classification-based analysis by providing insights into the functional role of different representations. See Appendix A for more detailed discussion of these interpretability approaches.
7 Conclusion
In this work, we examine how 16 transformer models from BERT to Llama 3.1 encode lexical identity and inflectional morphology. To do this, we train linear and non-linear classifiers on layer-wise activations to predict word lemmas and inflectional features, while using control tasks to measure classifier selectivity. Our investigation reveals consistent patterns across all models: lexical information concentrates in early layers and becomes increasingly non-linear in deeper layers, while inflectional information remains linearly accessible throughout the network. Moreover, our selectivity analysis shows that models primarily memorize lexical relationships, but learn generalizable patterns for inflection.
Crucially, these encoding patterns hold across different architectures, model sizes, and training procedures. These findings suggest that this linguistic information is essential for both next token prediction during pretraining and complex instruction-following, demonstrating that despite rapid advances in language technologies, certain aspects of how these models process language remain remarkably consistent.
8 Limitations
Representation Extraction for Decoder Models
Form and Function in English Inflection
English contains cases where different grammatical functions share the same surface form (e.g., infinitive and non-past verb forms). We do not explicitly look at these cases in our classification/probing experiments, but these ambiguities create opportunities to better examine how models separate form from function.
English-Only Analysis
Our analysis focuses solely on English, limiting generalization to languages with richer morphological systems, where the separation of lexical identity and inflectional morphology may differ in fundamental ways.
Indirect Nature of Classifiers
While our classifier methodology follows established best practices (Hewitt and Liang, 2019; Liu et al., 2019), we only detect correlations in hidden activations, not causal mechanisms. The increasing non-linearity of lexical encoding in deeper layers suggests that linear regression, MLP and random forest classifiers may not fully capture the complete picture of how information is organized.
Appendix A Additional Related Work
A.1 Advanced Probing Methodologies
Beyond standard linear probes, there are many sophisticated approaches to understand model representations. Amnesic probing (Elazar et al., 2021) removes specific information from representations to test whether it’s necessary for downstream tasks. Minimum description length probes (Voita and Titov, 2020) balance probe complexity with performance to avoid overfitting. Causal probing (Geiger et al., 2021) aims to establish causal rather than merely correlational relationships between representations and linguistic properties. Recently, Subramani et al. (2025) find that decoding from activations directly using the Logit Lens can be used to learn confidence estimators for tool-calling agents (nostalgebraist, 2020).
A.2 Model Manipulation and Steering
A.3 Feature Discovery and Mechanistic Interpretability
Appendix B Dataset Statistics
This section provides additional statistics and visualizations for the dataset used in our experiments. Only words containing alphabetic characters and apostrophes were considered. The data covers three main syntactic categories: nouns, verbs, and adjectives, with nouns being most frequent (49.5%). For our experimental setup, we split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels.
B.1 Category and Inflection Distribution
Table 3 shows the distribution of word categories in our dataset. Nouns constitute nearly half of all target words, followed by verbs and adjectives.

### v4

@@ 5.1 Results
Figure 4: Linguistic accuracy and classifier selectivity across model layers for English. The first two columns show lemma (top) and inflection (bottom) prediction accuracy using Linear Regression (left) and MLP (right) classifiers. The next two columns show classifier selectivity (difference between linguistic and control task accuracy) for the same tasks and classifiers. Higher selectivity indicates better generalization rather than memorization. Each line represents a different model. Multilingual results are shown in Figure˜5.
Figure 5: Cross-linguistic patterns in linguistic accuracy and classifier selectivity. The first two columns show lemma (top) and inflection (bottom) prediction accuracy using Linear Regression (left) and MLP (right) classifiers. The third column shows inflection prediction with Random Forest classifiers (lemma prediction is computationally infeasible due to the large number of classes). The rightmost columns show classifier selectivity for lemma (top) and inflection (bottom) tasks. Each line represents a different model-language combination.
We report results for lexical identity and inflectional morphology across classifier types and languages. The English set includes 19 models. We also evaluate five other languages using 6 multi/monolingual models. In addition to MLP and linear regression classifiers, we also use Random Forests for inflection (but not lemma, due to the large number of classes). Random Forest results are in Appendix˜I.
5.1.1 Lemma
Lemma task accuracy shows a sharp decline in accuracy with depth for linear regression classifiers, starting high (0.8–1.0) in the first layers and dropping in deeper layers (Figure˜4, top left). The size of the drop differs by model: DeBERTa-v3-Large falls from about 0.9 to 0.3, while Pythia-6.9B stays stronger later. By language, Turkish shows the largest decline (about 0.95 to 0.25), Russian maintains higher accuracy (about 0.6–0.8), and Chinese is in between. Two-layer MLPs raise accuracy and soften the decline (Figure˜4, top middle), with Pythia models strongest across layers and GPT-2, BERT, and Qwen2.5 gaining the most from added non-linearity. Gains are largest in Turkish, suggesting that in morphologically rich languages lemma information becomes more non-linear deeper in the network. Lemma selectivity is near zero across the board, suggesting that the probes mostly memorize.
5.1.2 Inflectional Morphology
Inflectional morphology task accuracy remains high (0.9–1.0) across layers and models for linear regression classifiers (Figure˜4, bottom left), with little change with depth across architectures and languages. English, German, French, and Russian are in the 0.90–1.0 range, while Turkish is lower (about 0.85–0.92), likely due to many-suffix combinations in agglutinative morphology. Comparative and superlative adjectives are consistently hardest. MLPs match linear probes for inflection (Figure˜4, bottom middle), suggesting inflection is linearly encoded. Random Forests are lower (0.65–0.90) and decline with depth, with Russian and German doing best and Turkish hardest (see Appendix˜I). Using control labels, inflection shows clear selectivity (about 0.4–0.6) across models and languages, with the highest values in Russian and German. Full curves appear in Figure˜4 and Figure˜5 and tables in Appendix˜J.
5.1.3 Probe Error Analysis
We also conduct an error analysis, grouping examples by part of speech and inflection type, and compare performance across frequency of occurrence. There is a clear correlation between frequency and accuracy: more frequent lemmas and inflections yield higher accuracy, particularly for lemmas, with rare words consistently the hardest. Among inflection types, comparative and superlative degrees, as well as low-count verb forms, exhibit the highest error rates across models. These trends are consistent across languages, with Turkish showing the strongest sensitivity to frequency. Detailed tables and plots are provided in Appendix˜J.
5.2 Analysis
Our experiments demonstrate clear patterns in how models process lexical identity and inflectional morphology. Here, we explore further and analyze six key aspects of these representations: cross-linguistic trends, representation structure, attention head outputs vs. residual stream, training dynamics, steering experiments, and intrinsic dimensionality.
Cross-Linguistic Trends
Across six languages, lemma is most readable early and declines with depth, while inflection stays high across layers (Figure˜5). Turkish shows the steepest lemma drop; Russian stays highest overall. These trends align with reports that agglutinative systems challenge current LMs.
Linear Representations
We compare linear probes with two-layer MLPs using the linear separability gap from §˜2. For inflection, gaps are near zero (sometimes slightly positive or negative), indicating that a linear map already captures most of the signal. For lemma, gaps are negative in early layers and shrink later, reflecting that regularized linear regression often outperforms a small MLP, especially where lexical information is concentrated and strongly tied to item frequency.
Intrinsic Dimensionality
To characterize how representations evolve through depth, we measure intrinsic dimensionality with PCA (Tables˜2 and 11). Models fall into two groups: some show gradual dimensionality decline (BERT, DeBERTa, Gemma, Llama 3.1), while others have middle layers where a single component explains most variance (GPT-2, Qwen2.5, Pythia, OLMo-2). Some families show gradual decline with depth; others have highly compressed mid-layers. Low intrinsic dimensionality has practical implications: we show below that steering vectors can more effectively manipulate model behavior by targeting compressed middle-layer representations.
Attention vs. Residual Stream
Comparing averaged attention head outputs to the residual stream, lemma information drops in attention outputs in middle layers while the residual stream stays higher. Inflection remains strong in both, with the residual stream slightly better in the middle. This supports the view that attention focuses on contextual relations, while MLP blocks store more parametric knowledge. See Appendix˜G.
Training Dynamics
We evaluate checkpoints for OLMo-2-7B and Pythia-6.9B. Three patterns emerge. First, the split between lemma and inflection appears early: lemma accuracy peaks at early steps and declines gradually with more training, while inflection accuracy is high and flat from the start. Second, the layer at which lemma is most readable shifts slightly toward earlier layers as training proceeds, consistent with increasing contextualization in middle and late layers. Third, these trends hold across both families despite different corpora and training scales, pointing to architecture-level behavior rather than a dataset artifact. Figures are in Figures˜24 and 25.
Steering Experiments
We apply steering vectors computed from class means for inflection categories and test several strengths. Steering is effective across most layers and models: mean probability change and flip rates are near 1.0 in many regions, showing that inflectional directions are functionally manipulable. The models which are least steerable are the same ones which have a low intrinsic dimensionality in mid-layers, suggesting that compressed representations are harder to adjust with simple vector shifts. See Appendix˜H for full details.
6 Related Work
Probing for linguistic information
Probing studies typically use supervised classifiers to predict linguistic properties from model representations (Alain & Bengio, 2017; Adi et al., 2017). Extensive work has established that early transformer models (BERT, GPT-2) learn hierarchical linguistic structures, with different layers specializing in different information types: lower layers capture surface features and morphology, middle layers encode syntax, and upper layers represent semantics and context (Jawahar et al., 2019; Tenney et al., 2019a; Rogers et al., 2020). More relevant to our work, Vulić et al. (2020) found that lexical information concentrates in lower layers, while Ethayarajh (2019) showed that representations become increasingly context-specific in higher layers.
Activation steering
Beyond probing, recent work has explored manipulating model behavior by intervening on internal representations. This includes steering vectors (Subramani et al., 2022), inference-time interventions (Li et al., 2023), representation editing (Meng et al., 2022), sparse autoencoders for feature discovery (Bricken et al., 2023), and causal mediation analysis (Vig et al., 2020). While these methods typically evaluate changes in model outputs, our steering experiments focus on measuring representational changes. See Appendix˜B for more detailed discussion.
7 Conclusion
In this paper, we examined 25 transformer models and found evidence that modern LMs rediscover the classical NLP pipeline: syntax is most accessible first, then semantic and entity-level information, and finally discourse. At the same time, we observe that larger contemporary models shift these peaks toward earlier layers, suggesting that stronger capacity and training lead to useful representations emerging sooner. Our in-depth study of lexical identity and inflectional morphology shows a clear pattern. Lemma information concentrates early and becomes increasingly non-linear with depth, while inflectional features remain linearly accessible throughout layers and across languages. Analyses of attention vs. residual streams, pretraining checkpoints, and steering vectors further locate where this information lives, when it appears, and how it can be manipulated. These results hold across architectures, sizes, and training regimes, pointing to robust representational regularities that persist despite rapid advances in model scale and training practice.
8 Reproducibility Statement
Our GitHub repository is available at https://github.com/ml5885/model_internal_sleuthing. It contains code to reproduce dataset construction, probing experiments, and all plots and analyses. The main paper specifies the probe design and metrics (§˜2), datasets and model suite (§˜3 and Table˜1), and evaluation summaries for the classical pipeline and for lexical identity and inflectional morphology (§§˜4 and 5). The appendix provides complete classifier and training details, dataset statistics, and full-resolution figure grids referenced in the text. Together, these materials are intended to enable end-to-end reproduction of our results.
Appendix A Limitations
Representation Extraction for Decoder Models
Form and Function in Inflection
Some languages contain cases where different grammatical functions share the same surface form (e.g., , infinitive and non-past verb forms in English). We do not explicitly examine these cases in our classification experiments, but these ambiguities create opportunities to better examine how models separate form from function across languages.
Indirect Nature of Classifiers
While our classifier methodology follows established best practices (Hewitt & Liang, 2019; Liu et al., 2019), we only detect correlations in hidden activations, not causal mechanisms.
Scope of Steering Experiments
Our steering vector experiments measure changes in classifier performance rather than downstream model outputs. Evaluating effects on actual model generation would require more complex experimental designs to control for confounding factors and ensure that observed changes result from the intended representational modifications rather than other influences.
Appendix B Additional Related Work
B.1 Advanced Probing Methodologies
Beyond standard linear probes, there are many sophisticated approaches to understand model representations. Amnesic probing (Elazar et al., 2021) removes specific information from representations to test whether it’s necessary for downstream tasks. Minimum description length probes (Voita & Titov, 2020) balance probe complexity with performance to avoid overfitting. Causal probing (Geiger et al., 2021) aims to establish causal rather than merely correlational relationships between representations and linguistic properties. Recently, Subramani et al. (2025) find that decoding from activations directly using the Logit Lens can be used to learn confidence estimators for tool-calling agents (nostalgebraist, 2020).
B.2 Model Manipulation and Steering
Recent work has also examined how multilingual models like mT5 and ByT5 encode morphological information differently across languages (Dang et al., 2024), finding that tokenization strategies significantly impact morphological representation quality, particularly for morphologically rich languages.
B.3 Feature Discovery and Mechanistic Interpretability
Appendix C Training Details
We stratify all datasets into train, validation and test splits. We train classifiers using the training set, select hyperparameters using the validation set, and evaluate on the held-out test set using accuracy and macro F1. For linear regression, we use ridge regularization with $\lambda=0.01$. For random forests, we tune the number of trees (50, 100, 200) and maximum depth (5, 10, 20, None) using grid search. For MLPs, we use a hidden layer size of 64, learning rate of 0.001, weight decay of 0.01, and train for up to 100 epochs with early stopping based on validation loss. We solve equation˜1 in closed-form to identify $W\in\mathbb{R}^{d\times c}$, and learn $W_{1}$ and $W_{2}$ for the MLP via stochastic gradient descent using the AdamW optimizer on cross-entropy loss (Kingma & Ba, 2014).   [^4] We use weight decay for regularization for MLP models.roximators, capable of approximating any continuous function to arbitrary precision given sufficient width (Hornik et al., 1989).
Appendix D Full Results for the Classical NLP Pipeline
Figure 6: Full heatmaps for MLP probe accuracy across all tasks, models, and layers. Rows show tasks; columns show models; each cell shows accuracy by layer depth. Early layers favor syntax, middle layers semantics, late layers discourse.
Figure 7: Full heatmaps for linear probe accuracy across all tasks, models, and layers. Trends mirror the MLP version but with stronger model-size effects in deeper layers.
Figure 8: Full heatmaps for MLP probe selectivity (real vs. control task accuracy). Higher values indicate true signal rather than memorization.
Figure 9: Full heatmaps for linear probe selectivity (real vs. control task accuracy). Selectivity is generally higher for semantic and discourse tasks.
Figure 10: Full Expected Layer and Center of Gravity plots with $\tau_{\text{lin}}$ and $\tau_{\text{MLP}}$ selectivity scores for all models. We observe early peaks for syntax, middle peaks for semantics, and late peaks for discourse across architectures and sizes.
Appendix E Additional Analysis
Figure 11: Intrinsic dimensionality across layers. Lines show fraction of PCA components needed to reach variance thresholds (50–100%). Models with strong mid-layer compression (few components for high variance) align with the inflection-is-linear, lemma-is-nonlinear split. Full curves per model appear in Figure˜12.
E.1 Intrinsic Dimensionality Results
@@
Figure 12: Intrinsic dimensionality curves for all models for English. Each subplot shows the relationship between the percentage of maximum PCA components (x-axis) and the percentage of explained variance (y-axis) across different layers. The color gradient from purple (early layers, 0%) to yellow (late layers, 100%) indicates the relative layer depth within each model. Models like BERT, Gemma, and Llama show similar compression patterns, while GPT-2 variants, Qwen and Pythia exhibit opposite trends in their middle layers.
Appendix F Dataset Statistics
This section provides statistics and visualizations for the datasets and models used in our experiments across all six languages. Only words containing alphabetic characters and apostrophes were considered.
Language Total Words Unique Lemmas Unique Forms Inflection Types Sentences Avg. Length
English 54,816 7,848 11,720 8 8,415 6.5
Chinese 44,166 11,184 11,237 4 7,892 5.8
German 84,710 24,140 31,890 9 9,234 7.3
French 115,847 13,804 24,485 6 8,765 6.6
Russian 193,320 20,943 59,830 8 10,234 7.1
Turkish 20,881 3,776 11,680 7 6,789 6.4
Table 3: Dataset statistics across all six languages. Russian has the largest dataset and the highest number of unique forms, reflecting its rich inflectional morphology. Turkish has the fewest total words and lemmas, while Chinese has the fewest inflection types.
F.1 English Dataset Details
For the English GUM corpus specifically, the data covers three main syntactic categories: nouns (49.5%), verbs (31.2%), and adjectives (19.4%).
Table˜4(a) shows the distribution of word categories in the English dataset, and Table˜4(b) presents the distribution of inflection categories.